OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2)

OpenStudy (anonymous):

The graph has 2 points Point A: (-2, 0) Point B: (0, -1)

OpenStudy (anonymous):

point a because perpendicular means the opposite so therefore the opposite of the y axis will be -2

OpenStudy (anonymous):

Y=2x+4 Y=-2x+4 Y=-1/2x+2 Y=2x-1

OpenStudy (anonymous):

@denishays111

Nnesha (nnesha):

first you need to find slope of two points (-2,0)(0,-1) formula to find slope is \[\huge\rm \frac{ y_2 -y_1}{ x_2 -x_1 }\] plug in y and x values

OpenStudy (anonymous):

(-1)-0= 0-(-2)

OpenStudy (anonymous):

So -1/2

Nnesha (nnesha):

yep right :=) now perpendicular slope is negative reciprocal for example if slope of first equation is - a/b then slope of perpendicular equation would be positive b/a

Nnesha (nnesha):

so what is the slope of 2nd equation ?

OpenStudy (anonymous):

-2x?

OpenStudy (anonymous):

Sorry 2x

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

So would the answer be y=2x+4

Nnesha (nnesha):

so equation of slope intercept form is y =mx+b where m is slope so substitute m for 2 and pick on of order pair to plug in value of x and y solve for b

Nnesha (nnesha):

so how did you get A ? and how do you know it's not D ?

OpenStudy (anonymous):

Because when i look at the line The y intercept must pass through (-1, 2)

OpenStudy (anonymous):

So i thought the line would be slanted the other way

OpenStudy (anonymous):

So its d?

Nnesha (nnesha):

perpendicular line passes through the point (-1,2) so substitute x and y for (-1,2)

Nnesha (nnesha):

nope first one is correct i just wanted to know how did YOU get A :=)

OpenStudy (anonymous):

Ok, thank you very much

Nnesha (nnesha):

my pleasure :=)